OpenStudy (anonymous):

Using complete sentences, explain how to graph y = one halfx - 3 using the slope-intercept method

OpenStudy (anonymous):

@Hero

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

ok so can u help me

OpenStudy (anonymous):

y = 1/2x - 3 -3 is your y intercept. to find you x intercept, sub in 0 for y... 0 = 1/2x - 3 -1/2x = -3 x = -3/(-1/2) x = -3 (-2) = 6 <-- your x intercept. your slope is 1/2....put a dot on the y axis at -3, go up 1 and over 2, then up 1 and over 2. It should look like this... http://www.quickmath.com/webMathematica3/quickmath/graphs/equations/basic.jsp#c=graph_stepsgraphequation&v1=y+%3D+1%2F2x+-+3

OpenStudy (anonymous):

thank u so much